OhSoPoor, Inc (OSP) was exceptionally short on cash during 2019 and issued a series of loans to cover their cash shortfall. The market rate of interest for all the bonds was 5%. For each bond, record the journal entry following transactions:

(a) Theissuanceofthebond.
(b) Thefirstinterestpaymentafterissuanceofthebond. (c) TheinterestpaymentoraccrualatDecember31,2019.

Bond A: This is a zero-interest, $100,000 bond which matures in 5 years and is issued on January 1, 2019.

Bond B: This bond is a 15-year, $500,000 bond which pays interest at 6%. This bond was issued on February 1, 2019 and pays interest on June 30, and December 31 of each year.

Bond C: This is a 10-year, 4%, $200,000 bond. This bond was issued on April 1, 2019 and pays interest on March 31 and September 30 of each year.

Bond D: This is a 7.5 year, 8% bond with a maturity value of $400,000. The bond pays interest at the end of every calendar quarter. This bond was issued on July 1, 2019.
Answer #1
If the market rate and the contract rate of interest are different, the bonds would sell at
premium or discount.
If the contract rate is less than market rate, the bonds sell at discount
If the contract rate is more than market rate, the bonds sell at premium
The formula to compute the issue price/selling price of Bond is as under
PV = r * B *( 1 - (1+mr) ^ -(n*p) + B/(1+mr)^(n*p)
where, PV = Issue price of Bond
B = Face value/Par Value of the Bond
r = coupon rate of the Bond
mr = market rate of interest
n = no of years to maturity
p = frequency of bond interest payment in a year
Computation of issue price of bonds by using the formula above
Bond A
= 0 * 100,000 * (1-(1.05) ^-(5*1))/r + 100,000/(1.05) ^ 5
= 0+ (100,000)/(1.2763)
= 0+ 78,351
Issue Price = $ 78,351
Bond B
= 0.06 * 500,000 * (1-((1.05) ^-(15*2))/.05 + 500,000/((1.05) ^ (15*2))
= 30,000 * 1-(1.05 ^ -30)/0.05 + 500,000/(1.05^30)
=30,000 * (1 - (0.2314)/0.05) + 500,000/4.3219
= 30,000 * 15.372 + 115690
= 461160+115690
= 576850
Issue Price = $ 576,850
Bond C
= 0.04 * 200,000 * (1-((1.05) ^-(10*2))/.05 + 200,000/((1.05) ^ (10*2))
= 8000 * (1-0.3769)/0.05 + 200000/2.6533
= 8000 * 0.6231/0.05 + 75378
= 8000 * 12.462 + 75378
= 99696 + 75378
= 175074
Issue Price = $ 175,074
Bond D
= 0.08 * 400,000 * (1-((1.05) ^-(7.5*4))/.05 + 400,000/((1.05) ^ (7.5*4))
= 0.08 * 400,000 * (1-((1.05) ^-30)/.05 + 400,000/((1.05) ^ (30))
= 32000 * (1-0.2314)/.05 + 400000/4.3219
= 32000 * 0.7686/.05 +92552
=32000 * 15.372 +92552
= 584456 + 92552
= 677008
Issue Price = $ 677,008
Journal Entries
Date Account Title and Explanation Debit Credit
Jan 1 Cash        78,351
Discount on Bond A payable        21,649
Bond A payable       100,000
(to record the issue of issue of zero interest Bond A with par value of $ 100,000 at discount)
Feb 1 Cash       576,850
Premium on Bond B payable        76,850
Bond B payable       500,000
(to record the issue of issue of 6% Bond B with par value of $ 500,000 at premium)
Apr 1 Cash       175,074
Discount on Bond C payable        24,926
Bond C payable       200,000
(to record the issue of issue of 4% Bond C with par value of $ 200,000 at discount)
Jul 1 Cash       677,008
Premium on Bond D payable       277,008
Bond D payable       400,000
(to record the issue of issue of 6% Bond B with par value of $ 500,000 at premium)
June 30 Interest Expense        10,365
Premium on Bond B payable          2,135
Cash        12,500
(to record the interest expense and amortization of premium payable on Bond B)
Sep 30 Interest Expense        11,739
Discount on Bond C payable          3,739
Cash          8,000
(to record the interest expense and amortization of discount on Bond C)
Sep 30 Interest Expense          1,233
Premium on Bond D payable          9,233
Cash          8,000
(to record the interest expense and amortization of premium payable on Bond D)
Dec 31 Interest Expense        12,438
Premium on Bond B payable          2,562
Cash        12,500
(to record the interest expense and amortization of premium payable on Bond B)
Dec 31 Interest Expense          4,330
Discount on Bond A payable          4,330
(to record the interest expense and amortization of discount on Bond A)
Dec 31 Interest Expenses          3,869
Discount on Bond C payable          1,869
Accrued Expenses          2,000
(to record the interest expense and amortization of discount on Bond C)
Dec 31 Interest Expense          1,233
Premium on Bond D payable          9,233
Cash          8,000
(to record the interest expense and amortization of premium payable on Bond D)